# Recov-Vision: Linking Street View Imagery and Vision-Language Models for Post-Disaster Recovery

### v1

@@ Abstract
Building-level occupancy after disasters is vital for triage, inspections, utility re‑energization, and equitable resource allocation. Overhead imagery provides rapid coverage but often misses facade and access cues that determine habitability, while street‑view imagery captures those details but is sparse and difficult to align with parcels. We present FacadeTrack, a street‑level, language‑guided framework that links panoramic video to parcels, rectifies views to facades, and elicits interpretable attributes (for example, entry blockage, temporary coverings, localized debris) that drive two decision strategies: a transparent one‑stage rule and a two‑stage design that separates perception from conservative reasoning. Evaluated across two post–Hurricane Helene surveys, the two‑stage approach achieves precision 0.927, recall 0.781, and F‑1 score 0.848, compared with the one‑stage baseline at precision 0.943, recall 0.728, and F‑1 score 0.822. Beyond accuracy, intermediate attributes and spatial diagnostics reveal where and why residual errors occur, enabling targeted quality control. The pipeline provides auditable, scalable occupancy assessments suitable for integration into geospatial and emergency‑management workflows.
I Introduction
I-A Motivation & problem statement
@@
Rapid, building-level occupancy status after a disaster is a time-critical input for triage, utility restoration, inspections, temporary shelter planning, and equitable resource allocation. Overhead (nadir) imagery arrives quickly and covers large areas, but its roof-dominant perspective often overlooks the very cues that determine whether people can safely inhabit a building–open or blocked entries, temporary coverings, localized debris, and ad-hoc repairs. Even state-of-the-art post-disaster overhead methods struggle when habitability hinges on facade and access conditions rather than roof damage alone [1, 2, 3]. Street-based visual surveys and door-to-door assessments supply those missing signals, yet they are inherently sparse and viewpoint-dependent, constrained to the road network and acquisition schedules, and their alignment with parcel inventories is noisy in practice [4, 5, 6, 7]. As a result, analysts often face a slow, manual reconciliation of overhead and ground evidence that is difficult to reproduce or audit at scale, and that typically occurs only once immediately after impact rather than being repeated to track recovery over time.
@@
Recent progress in vision-language models (VLMs) could provide a mechanism for turning heterogeneous visual evidence into operational decisions without hand-crafted taxonomies for every new event. By combining open-vocabulary recognition, instruction following, and cross-modal reasoning, VLMs can be prompted to connect small visual cues with higher-level statements about building condition and use [8]. Early domain examples point in this direction—for instance, explanation-driven change reasoning on paired remote-sensing images [9], multi-hazard benchmarks for multimodal analysis [10], and large-model-assisted enrichment of facade attributes [11]. These studies establish feasibility and hint at auditability through natural-language rationales, but they stop short of a reproducible, street-level pipeline centered on post-disaster occupancy inference.
@@
A key limitation of past works is stitching together complementary evidence across views, turning ambiguous signals into explicit intermediate attributes, and proving reliability against specialized baselines. Overhead and street-level perspectives capture different failure modes–roof versus entry and facade–and simple late fusion rarely suffices when coverage is uneven or views are occluded [5]. Small, indirect indicators such as vehicle mix, tarps behind foliage, or shadows around doorways need to be promoted into structured, inspectable attributes so that an assessor can see how the model moved from evidence to judgment [11, 8]. And because domain-tuned aerial and street-view methods already achieve high accuracy–often in the face of severe class imbalance–claims of progress require statistical comparisons, spatial localization of errors, and analysis of failure cases tied to viewpoint and occlusion [3, 1, 2, 4]. These gaps become more pressing each year as large-scale flood and storm disasters affect growing coastal and urban populations, and as agencies need scalable, auditable assessments to guide and defend resource-allocation decisions [12].
Agencies need parcel-level occupancy within days to triage inspections, re-energize utilities, plan temporary shelter, and allocate resources equitably. Yet nadir imagery systematically misses facade and access cues—blocked entries, emergency placards, tarps, exterior mud—that determine habitability, while ad-hoc reconciliation of ground photos is slow and hard to audit. The study’s rationale is to close this gap with a reproducible, street-level, language-guided pipeline that makes evidence-to-judgment traceable: it links 5.6K 360° video to parcels, estimates heading from short GPS windows, and uses a simple yaw formulation to generate facade-centered rectified views. It then elicits nine interpretable visual attributes and applies two complementary decision strategies—a transparent one-stage rule with vehicle adjustment and a two-stage design that separates perception from conservative reasoning—so operating policies can be tuned without retraining. Together, these choices address the core needs of auditability, scalability, and defensibility for post-disaster occupancy decisions.
I-B Research objectives and significance
@@
This work seeks to develop a street-level, language-guided framework for post-disaster building occupancy assessment that combines street-level imaging, automated facade rectification, and vision-language reasoning to produce parcel-level occupancy labels and recovery trajectories. Our objective is to develop a reproducible street-level pipeline that combines facade rectification, VLM attribute extraction, and conservative reasoning to monitor parcel-level occupancy change across repeat visits. This objective set emphasizes operational reproducibility, interpretable intermediate evidence, and statistically grounded temporal and spatial evaluation for recovery monitoring.
[…]
I-C Contributions
Despite major progress in overhead change detection, habitability remains driven by facade and access conditions that nadir views routinely miss—blocked entries, emergency placards, exterior mud, ad-hoc repairs—forcing analysts into slow, ad-hoc reconciliation of roof-dominant evidence with scattered ground photos that are difficult to reproduce or audit. Street-view imagery contains the missing cues but is sparse, viewpoint-dependent, and noisy to align with parcel inventories in practice, while recent VLM efforts demonstrate feasibility without delivering a reproducible, street-level pipeline for post-disaster occupancy inference. What remains unsolved is threefold: cross-view integration when coverage is uneven or occluded, promoting weak, ambiguous signals (such as vehicles behind foliage or mud lines) into structured, inspectable attributes, and rigorous validation against specialized baselines with statistical tests, spatial localization of residual error, and failure-mode analysis under class imbalance. These persistent knowledge gaps motivate a different starting point.
Operationally, agencies need parcel-level occupancy within days to triage inspections and re-energize utilities, but current workflows rely on overhead-only triage or manual, non-auditable fusion of ground evidence, with no standard way to trace evidence to attributes to decision, tune risk-aware thresholds without retraining, report temporal change (recovered/deteriorated) credibly, or direct QA/QC using spatial error diagnostics. In short, practice lacks an end-to-end, repeatable capture-to-label chain—from panoramic video and GPS linkage through facade rectification to interpretable attributes and defensible decisions—that can scale while preserving auditability and privacy. Establishing this tooling and evaluation scaffolding is the point of departure for the present study: to replace ad-hoc reconciliation with a modular, language-guided, street-level pipeline explicitly designed for measurement fidelity, policy tuning, and targeted human oversight.
[…]
I-D Paper organization
@@
Section II reviews (i) vision-language foundations, (ii) aerial/satellite imagery-based damage assessment, and (iii) street view-related pipelines. Section III introduces data collection, frame-parcel linkage, panoramic rectification, attribute labeling, and the two decision strategies. Section IV reports classification performance, representative examples, two-visit occupancy change, and spatial error patterns. Section V discusses deployment implications, limitations, and future directions. Section VI concludes with key contributions and implications for rapid, auditable occupancy monitoring.
II Related work
II-A Aerial and satellite imagery for post-disaster damage
@@
Deep learning with overhead imagery has enabled rapid, wide-area mapping of building damage through semantic segmentation and change detection. Synthesizing this progress, [13] survey architectures, datasets, and open challenges for remote-sensing-based damage detection. Early pipelines used convolutional models on open aerial data to map impacts [14], while operational analyses documented how nadir views are confounded by cloud, canopy, and shadows, which obscure key cues and complicate severity estimation [15]. Alongside dataset curation, typology-driven formulations argue for more actionable, interpretable categories aligned with decision needs rather than coarse severity [16].
Modeling advances improved temporal fusion and representation capacity. Hierarchical transformers scale to large areas while separating fine-grained categories [2]. For example, Siamixformer introduces a fully transformer-based Siamese architecture for bi-temporal change reasoning [17] and BDANet combines multiscale features with cross-directional attention to capture both local and global cues [1]. UAV and oblique-aerial studies extend beyond satellites to quantify post-event damage with tailored models and views [18, 19, 20].
Overall, overhead approaches deliver speed, coverage, and mature change-detection tool chains, and remain indispensable for regional triage. However, they struggle when habitability hinges on entries and facades rather than roof conditions [21], and performance can degrade under severe class imbalance or subtle states [22, 15]. These limitations motivate complementing the nadir evidence with street-level cues and adopting reasoning pipelines that can articulate functional implications at the building scale.
@@
II-B Street-view imagery for urban condition and post-event assessment
Street-view imagery (SVI) provides facade-level detail that reflects access and near-term functionality. A panoramic SVI survey documents broad applications in data-driven urban research and the methodological considerations that come with roadside acquisition [6], while open-source pipelines have been introduced to streamline ingest, filtering, and sampling for research and operations [7]. In post-disaster settings, bi-temporal SVI supports hyperlocal change assessment at the building scale [4]; multi-modal fusion with structured attributes improves granularity and predictive performance for damage outcomes [5].
Coverage and viewpoint variability motivate multi-view designs and fusion with other sensors. Leveraging multiple SVI vantage points improves robustness under occlusions and oblique angles [23], and fusing SVI with remote sensing supports broader risk and exposure modeling in flood contexts [24, 25]. These advances show how ground imagery complements overhead views with human-interpretable evidence, even as road-network constraints and capture schedules limit universal availability.
II-C Vision-language and multimodal foundations for the built environment and disasters
Vision-language models offer natural-language interfaces to visual evidence, enabling open-vocabulary recognition and cross-modal reasoning with limited task-specific training. [8] survey emerging applications in remote sensing; recent work tailors VLMs to disaster analysis and built-environment tasks, enabling models to capture context-specific features such as flood extent, building accessibility, and structural condition that are essential for recovery monitoring. ChangeChat demonstrates explanation-driven change reasoning on paired remote-sensing images [9], and DisasterM3 provides a multi-hazard, multi-sensor benchmark to evaluate multimodal models under diverse conditions [10]. In street-level contexts, language-conditioned estimation of lowest-floor elevation illustrates how VLMs can extract actionable attributes from facade imagery [26], and large-model-assisted enrichment frameworks expand architectural attributes at scale using SVI [11]. Complementary efforts on knowledge fusion and data generation aim to structure damage information and mitigate imbalance by combining imagery with expert priors within VLM-driven pipelines [27].
Taken together, this survey of past studies suggests that language-guided reasoning can connect small, ambiguous cues to decision-ready statements while exposing intermediate rationales for auditing. Prompt sensitivity, small-object perception limits, and the need for domain-native resources remain practical hurdles, but as coastal and urban risks intensify [12], the combination of overhead imaging, SVI, and VLMs offers a promising basis for auditable, building-level occupancy inference. Our work builds on these strengths while addressing the remaining gaps around cross-view integration, transparent attribute-to-decision tracing, and rigorous evaluation against specialized baselines.
III Methodology
@@ III-A Data collection
As a test case, we collected imagery from two field visits to areas impacted by Hurricane Helene. Helene made landfall as a Category 4 hurricane in the Big Bend region of Florida on September 26, 2024, before traveling inland and stalling over the North Carolina/Tennessee border on September 27–28. In total, Helene is estimated to have generated more than \$78 billion in economic losses [28] and at least 250 fatalities. Of these, 95 were directly attributable to freshwater flooding (which includes landslides and debris flows) in North Carolina, South Carolina, and Tennessee, making it the deadliest hurricane to make landfall in the continental United States since Hurricane Katrina (2005) [29]. Notably, Helene’s heavy precipitation (i.e., 20–30 inches, corresponding to a 1000-year rainfall event in western North Carolina) over a region characterized by steep topography generated widespread flooding, damaging tens of thousands of homes and cutting off entire communities from access to critical facilities such as water, sewer, electricity, telecommunications, and healthcare [30]. The southern Appalachian Mountains exhibit particularly low flood insurance penetration (i.e., only 1.5% of buildings had coverage [31]), raising questions about how quickly the region will recover and what household recovery will look like.
@@
Our study focused on several flooded neighborhoods in Swannanoa and Black Mountain in Buncombe County, North Carolina. (Swannanoa is a census-designated place located between the City of Asheville and the Town of Black Mountain.) Both communities are located along the Swannanoa River, which is a tributary of the French Broad River. The key survey areas are shown in Figure 1. Prior post-disaster surveys (e.g., Red Cross assessments and county emergency management reports [32]) identified these neighborhoods as among the most heavily affected communities in the region. Our field reconnaissance in December 2024 confirmed the high density of visible damage, making them appropriate testbeds for property-level monitoring. The damaged areas comprised a mix of suburban and semi-rural parcels and had relatively accessible road networks, enabling repeated street-level imaging to support temporal analysis of recovery. In addition to the initial reconnaissance visit conducted in December 2024, we collected street-level data across two field campaigns in March 2025 and June 2025.
@@
Figure 1: Study area in broader Asheville area, North Carolina, showing residential parcels and drive routes used for panoramic data collection.
@@
During each field visit, we collected street‐level, 360° imagery using a GoPro Max panoramic camera mounted on a rigid roof–rack mast (approximately 2 m above ground level) on a passenger vehicle. Videos were recorded at the camera’s native panoramic resolution (5.6K) at 30 fps to support frame‐level alignment with GPS. Each recording was accompanied by a synchronized per‐video GPS log (CSV) capturing latitude/longitude, speed, and course, enabling subsequent geospatial matching of frames to reference parcels (Section III-B). The camera was leveled and oriented with the vehicle’s leftward heading to better capture the building facades, with horizon lock enabled to stabilize the panoramic stream. For each campaign, operators maintained a simple time‐stamped field log noting start/stop points, detours, obstructions, and environmental conditions to facilitate post hoc quality control and alignment. All data were captured from public rights‐of‐way; no entry onto private property occurred. To protect privacy, personally identifying details (e.g., faces and license plates) were blurred or removed prior to the dissemination of imagery.
III-B Data pre-processing

### v2

@@ Abstract
Building-level occupancy after disasters is critical for triage, inspections, utility re-energization, and equitable resource allocation. Overhead imagery provides rapid coverage but often misses facade and access cues that determine habitability, while street-view imagery captures those details but is sparse and hard to align with parcels. We present Recov-Vision, a street-level, language-guided framework that links panoramic video to parcels, rectifies views to facades, and elicits interpretable attributes (e.g., entry blockage, temporary coverings, localized debris) that drive two decision strategies: a transparent one-stage rule and a two-stage design separating perception from conservative reasoning. Evaluated across two post-Hurricane Helene surveys, the reasoning approach attains higher recall with comparable overall agreement to a one-stage baseline and reproduces the ground-truth net recovery (+14), while intermediate attributes and spatial diagnostics expose where and why residual errors occur. The pipeline provides auditable, scalable occupancy assessments suitable for integration into geospatial and emergency-management workflows.
keywords:
Street-view imagery, Vision-language models (VLMs), Geospatial AI, Post-disaster occupancy assessment, Parcel-scale recovery monitoring, Resilience analytics
1 Introduction
1.1 Motivation & problem statement
@@
Rapid, building-level occupancy status after a disaster is a time-critical input for triage, utility restoration, inspections, temporary shelter planning, and equitable resource allocation. Overhead (nadir) imagery arrives quickly and covers large areas, but its roof-dominant perspective often overlooks the very cues that determine whether people can safely inhabit a building—open or blocked entries, temporary coverings, localized debris, and ad-hoc repairs. Even state-of-the-art post-disaster overhead methods struggle when habitability hinges on facade and access conditions rather than roof damage alone [shen_bdanet_2022, kaur_largescale_2023, rahnemoonfar_floodnet_2021].
Street-based visual surveys and door-to-door assessments supply those missing signals, yet they are inherently sparse and viewpoint-dependent. They are constrained to the road network and acquisition schedules, and their alignment with parcel inventories is often noisy [yang_hyperlocal_2025, xue_post-hurricane_2024, cinnamon_panoramic_2021, martell_open-source_2024]. As a result, analysts face a slow, manual reconciliation of overhead and ground evidence that is difficult to reproduce or audit at scale. This process typically occurs only once immediately after impact, rather than being repeated to track recovery over time.
@@
Recent progress in vision-language models (VLMs) could provide a mechanism for turning heterogeneous visual evidence into operational decisions without hand-crafted taxonomies for every new event. By combining open-vocabulary recognition, instruction following, and cross-modal reasoning, VLMs can connect small visual cues with higher-level statements about building condition and use [weng_vision-language_2025]. While early examples establish feasibility—ranging from change reasoning [deng_changechat_2025] to language-conditioned elevation estimation [ho_integrated_2025]—they stop short of a reproducible, street-level pipeline centered on occupancy inference.
@@
A key limitation of past work is the inability to stitch together complementary evidence across views to turn ambiguous signals into explicit intermediate attributes. Small indicators such as vehicle mix, tarps behind foliage, or shadows around doorways need to be promoted into structured, inspectable attributes so that an assessor can verify how the model moved from evidence to judgment [liang_openfacades_2025]. Furthermore, because domain-tuned aerial and street-view methods already achieve high accuracy, claims of progress require statistical comparisons, spatial localization of errors, and analysis of failure cases tied to viewpoint and occlusion. These gaps become more pressing as agencies seek scalable, auditable assessments to defend resource-allocation decisions [paulik_local_2025].
1.2 Research objectives and significance
@@
This work develops a street-level, language-guided framework for post-disaster building occupancy assessment. We combine panoramic imaging, automated facade rectification, and vision-language reasoning to produce parcel-level occupancy labels and recovery trajectories.
The primary objective is to create a reproducible pipeline that transforms raw street-level video into auditable occupancy decisions. Unlike ”black box” end-to-end classifiers, our approach emphasizes intermediate evidence (facade rectification, attribute extraction) and conservative reasoning to monitor parcel-level change across repeat visits.
The significance of this study lies in two areas: measurement fidelity and operational targeting. By employing a two-stage reasoning strategy, the pipeline improves recall and accurately reproduces ground-truth net recovery rates, whereas standard baselines tend to overstate recovery. This yields credible metrics that decision-makers can act on. Spatial diagnostics reveal significant clustering in occupancy states and residual errors. By generating accuracy maps, the framework allows agencies to pinpoint pockets for targeted quality assurance and quality control (QA/QC) and field tasking rather than relying on diffuse, random reviews.
1.3 Contributions
To address the gaps in cross-view integration, attribute interpretability, and rigorous validation, this paper makes four main contributions:
  1. 1.
    A reproducible street-level capture-to-label pipeline. We document an end-to-end workflow that links 5.6K panoramic videos to parcels, estimates heading from GPS windows, and uses a simple yaw formulation to rectify frames into facade-centered views. These geometry and processing steps are fully specified to enable replication and auditing.
  2. 2.
    Interpretable, policy-tunable occupancy inference. We compare two decision designs using a shared vision extractor: a transparent one-stage rule and a two-stage perception-to-reasoning approach. By separating perception from logic, thresholds can be tuned for conservatism without retraining the vision component.
  3. 3.
    Temporally and spatially aware evaluation. We introduce a change framework (Recovered, Deteriorated, Stable) to quantify net recovery across campaigns. We utilize global Moran’s I to diagnose the geographic concentration of errors, enabling spatially targeted quality control.
  4. 4.
    Statistically grounded validation. We report paired bootstrap confidence intervals and McNemar tests to rigorously compare strategies. Our error taxonomy shows that the two-stage approach reduces ambiguity, guiding reviewers to a small, well-defined subset of difficult cases.
1.4 Paper organization
@@
Section 2 reviews vision-language foundations and existing damage assessment pipelines. Section 3 details the data collection, frame-parcel linkage, and decision strategies. Section 4 reports classification performance, recovery metrics, and spatial error patterns. Section 5 discusses deployment implications and limitations, followed by conclusions in Section 6.
2 Related work
2.1 Aerial and satellite imagery for post-disaster damage
@@
Deep learning with overhead imagery has become the standard for rapid, wide-area damage mapping. Synthesizing this progress, wang_deep_2024 survey architectures and open challenges, noting a shift from simple classification to semantic segmentation. Early pipelines successfully used convolutional models on open aerial data to map impacts [gupta_deep_2021], yet operational analyses have consistently documented how nadir views are confounded by environmental factors—clouds, canopy, and shadows—that obscure key cues and complicate severity estimation [qing_operational_2022]. Furthermore, the lack of large, balanced datasets remains a persistent obstacle [wang_building_2022]. While curated benchmarks such as FloodNet [rahnemoonfar_floodnet_2021] have standardized protocols, their scale is limited. To address the gap between raw detection and decision-making, typology-driven formulations have recently argued for actionable categories aligned with civil recovery needs rather than coarse severity scores [xiao_damagecat_2025].
Recent modeling advances have focused on improving representation capacity and temporal reasoning. Hierarchical transformers have emerged as a solution for scaling to large areas while preserving fine-grained detail [kaur_largescale_2023]. In the domain of change detection, Siamixformer introduced a fully transformer-based Siamese architecture for bi-temporal reasoning [mohammadian_siamixformer_2023], while BDANet combined multiscale features with cross-directional attention to better capture local and global context [shen_bdanet_2022]. Beyond single architectures, ensemble learning has been explored to enhance robustness across diverse hydrologic and geographic terrains [roohi_developing_2024].
Research has also expanded into end-to-end mapping systems that integrate imagery with GIS layers [braik_automated_2024] and urban remote-sensing workflows that pair deep learning with post-classification correction [techapinyawat_integrated_2024]. Recognizing the limits of satellite resolution, UAV and oblique-aerial studies have developed tailored models to quantify damage from lower-altitude perspectives [calantropio_deep_2021, gong_deep_2021, khankeshizadeh_novel_2024]. Some efforts have even coupled UAV data with GIS-based information systems for facade assessment [chen_gis-based_2023].
Despite these advances, overhead approaches remain fundamentally limited by their viewing angle. They struggle when habitability hinges on facade and access cues—such as boarded windows or blocked doors—rather than roof integrity [ho_flood-damagesense_2025], necessitating complementary ground-level evidence.
@@
2.2 Street-view imagery for urban condition and post-event assessment
Street-view imagery (SVI) provides the high-resolution, facade-level detail required for occupancy analysis. cinnamon_panoramic_2021 review the methodological considerations of roadside acquisition, while open-source pipelines have been introduced to streamline the ingest, filtering, and sampling of SVI for research [martell_open-source_2024]. Outside of disaster contexts, SVI has powered city-scale analyses of facade attributes [zhong_city-scale_2021], pavement defect mapping [maniat_deep_2021, kong_automatic_2022, wang_automatic_2021], and the identification of social vulnerability indicators [ogawa_deep_2023].
In disaster response, bi-temporal SVI has been applied to hyperlocal change assessment [yang_hyperlocal_2025], and multi-modal fusion studies have shown that combining SVI with structured attributes improves damage prediction accuracy [xue_post-hurricane_2024]. To address occlusion and viewing angle challenges, recent work has leveraged multiple SVI vantage points to improve robustness [gu_multiview_2025], or fused SVI with remote sensing for broader flood risk modeling [chen_multi-modal_2022, xing_flood_2023]. Operational guidance on reconnaissance-oriented SVI collection—covering route design and coverage—has also helped bridge the gap between research and deployment [errett_street_nodate].
While SVI captures the necessary visual data, most existing methods rely on ”black box” classifiers or simple late fusion. They lack the structured, transparent reasoning required to defend resource allocation decisions, particularly when views are occluded or oblique.
2.3 Vision-language and multimodal foundations
Vision-language models (VLMs) offer a mechanism to process visual evidence using open-vocabulary recognition and cross-modal reasoning. weng_vision-language_2025 survey emerging applications in remote sensing, noting the potential for VLMs to capture context-specific features like accessibility and structural condition. deng_changechat_2025 demonstrated explanation-driven change reasoning on paired images, while wang_disasterm3_2025 introduced a multi-hazard, multi-sensor benchmark to evaluate these models under diverse conditions.
Recent studies have begun to tailor VLMs to specific built-environment tasks. ho_integrated_2025 used language-conditioned estimation to determine lowest-floor elevation from street views, and liang_openfacades_2025 utilized large models to enrich facade attribute datasets at scale. In retrieval tasks, cross-view geo-localization has been used to surface relevant ground photos for overhead scenes [sogi_disaster_2024]. Beyond floods, multimodal systems have supported wildfire triage and infrastructure description [esparza_automated_2025, zhang_vision-enhanced_2025]. Complementary efforts on knowledge fusion aim to mitigate data scarcity by combining imagery with expert priors [wei_effective_nodate].
These studies establish the feasibility of VLM-based analysis but generally stop at detection or description. There remains a lack of integrated pipelines that use VLMs to drive a rigorous, reproducible, and auditable occupancy inference system capable of tracking recovery trajectories over time.
3 Methodology
@@ 3.1 Data collection
As a test case, we collected imagery from two field visits to areas impacted by Hurricane Helene. Helene made landfall as a Category 4 hurricane in the Big Bend region of Florida on September 26, 2024, before traveling inland and stalling over the North Carolina/Tennessee border on September 27–28. In total, Helene is estimated to have generated more than \$78 billion in economic losses [re2024natural] and at least 250 fatalities. Of these, 95 were directly attributable to freshwater flooding (which includes landslides and debris flows) in North Carolina, South Carolina, and Tennessee, making it the deadliest hurricane to make landfall in the continental United States since Hurricane Katrina (2005) [papin_climatology_2017]. Notably, Helene’s heavy precipitation (i.e., 20–30 inches, corresponding to a 1000-year rainfall event in western North Carolina) over a region characterized by steep topography generated widespread flooding, damaging tens of thousands of homes and cutting off entire communities from access to critical facilities such as water, sewer, electricity, telecommunications, and healthcare [re_natural_2024]. The southern Appalachian Mountains exhibit particularly low flood insurance penetration (i.e., only 1.5% of buildings had coverage [amorim_flooding_2025]), raising questions about how quickly the region will recover and what household recovery will look like.
@@
Our study focused on several flooded neighborhoods in Swannanoa and Black Mountain in Buncombe County, North Carolina. (Swannanoa is a census-designated place located between the City of Asheville and the Town of Black Mountain.) Both communities are located along the Swannanoa River, which is a tributary of the French Broad River. The key survey areas are shown in Figure 1. Prior post-disaster surveys (e.g., Red Cross assessments and county emergency management reports [Esri_COVID_Dashboard]) identified these neighborhoods as among the most heavily affected communities in the region. Our field reconnaissance in December 2024 confirmed the high density of visible damage, making them appropriate testbeds for property-level monitoring. The damaged areas comprised a mix of suburban and semi-rural parcels and had relatively accessible road networks, enabling repeated street-level imaging to support temporal analysis of recovery. In addition to the initial reconnaissance visit conducted in December 2024, we collected street-level data across two field campaigns in March 2025 and June 2025.
@@
Figure 1: Study area in broader Asheville area, North Carolina, showing residential parcels and drive routes used for panoramic data collection. Dots indicate buildings and lines indicate drive routes. Orange indicates routes/buildings from Visit 1, Green Visit 2, and Purple from both.
@@
During each field visit, we collected street‐level, 360° imagery using a GoPro Max panoramic camera mounted on a rigid roof-rack mast (approximately 2 m above ground level) on a passenger vehicle. Videos were recorded at the camera’s native panoramic resolution (5.6K) at 30 fps to support frame‐level alignment with GPS. Each recording was accompanied by a synchronized per‐video GPS log (CSV) capturing latitude/longitude, speed, and course, enabling subsequent geospatial matching of frames to reference parcels (Section 3.2). The camera was leveled and oriented with the vehicle’s leftward heading to better capture the building facades, with horizon lock enabled to stabilize the panoramic stream. For each campaign, operators maintained a simple time‐stamped field log noting start/stop points, detours, obstructions, and environmental conditions to facilitate post hoc quality control and alignment. All data were captured from public rights‐of‐way; no entry onto private property occurred. To protect privacy, personally identifying details (e.g., faces and license plates) were blurred or removed prior to the dissemination of imagery.
3.2 Data pre-processing